OpenStudy (anonymous):

Can someone please help with these geometry questions? There's no graph or anything, just questions. 1. Write the equation of the mirror line for the segment with endpoints (5, 4) and (-1, -1). 2. Find the reflection of the point A(3, -2) through the mirror line of y = 1/2x − 1.

OpenStudy (anonymous):

Please... I'm not asking for much..

OpenStudy (anonymous):

well can you write the equation of the line between the endpoints given in problem 1.

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

She said there is no graph so all we need to do here is get the slope-intercept equation

OpenStudy (anonymous):

For number 1?

OpenStudy (anonymous):

Yup I'm confident about it.....but what did you think we can do here?

OpenStudy (anonymous):

I was looking at number 2 first because number 1 is really easy.

OpenStudy (anonymous):

Lol I thought you're talking about #1. But ok let me work it out

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Oh this is simple all we have to look for first is a slope that is perpendicular to 1/2 Which is 2. So now we have to use the eq y-y1=m(x-x1) We know m is 2 and y1 = -2 and x1 = 3

OpenStudy (anonymous):

Plug in the given values into the equation. Can you solve for that @HallieBella?

OpenStudy (anonymous):

What's y in y-y1=m(x-x1) ? I see y1 = -2..

OpenStudy (anonymous):

This will help you with #1 if you do not get it http://www.purplemath.com/modules/strtlneq.htm If you get an answer for it, let us see and we'll see if you did right unless you know how to do this one already.

OpenStudy (anonymous):

F(x)=y you are solving for the function y

OpenStudy (anonymous):

I don't know.. I tried 2-(-2) = 4 then 2*(4-3) = 2 doesnt make sense.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Ok I'm going to show you the steps Y-y1=m(x-x1) Y-(-2)=-2(x-3) Distribute the -2 to x and -3 and you get Y+2=-2x+6 Subtract 2 on both sides and you get Y=-2x+4

OpenStudy (anonymous):

You understand

OpenStudy (anonymous):

Oh okay. I'm sorry. Math isn't my strongest point... So what would that make #1??

OpenStudy (anonymous):

ok so you know how to solve for the slope when you are given two coordinate points?

OpenStudy (anonymous):

m is the slope so \[m=\frac{ (y2-y1) }{ (x2-x1) }=\frac{ (4+1) }{ (5+1) }=\frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

now use the equation used to solve for y \[y-y1=m(x-x1)\] and solve for y

OpenStudy (anonymous):

so you get \[y-(-1)=\frac{ 5 }{ 6 }(x-(-1))\] then you have \[y+1=\frac{ 5 }{ 6 }(x+1)\] so you distribute the 5/6 and you get \[y+1=\frac{ 5 }{ 6 }x+\frac{ 5 }{ 6 }\] subtract 1 from both sides and you end up with \[y=\frac{ 5 }{ 6 }x-\frac{ 1 }{ 6 }\] your answer for the 1st question

OpenStudy (anonymous):

Thank you so much♥

OpenStudy (anonymous):

You welcome!!